OpenStudy (anonymous):

Find two vectors v1 and v2 and whose sum is(3,0,2) , where v1 is parallel to(-4,-2,-4) while v2 is perpendicular to (-4,-2,-4)

OpenStudy (anonymous):

suppose v1=(a1, a2, a3) and v2=(b1, b2, b3) a1 + b1= 3 a2 + b2= 0 a3 + b3=2 a1 = -4k a2= -2k a3= -4k and -4 b1- 2b2 - 4b3=0

OpenStudy (anonymous):

a1+a2+a3+b1+b2+b3=5 (-4k-2k-4k) + (b1+b2+b3)=5 b1+b2+b3=5 +10k and -4 b1- 2b2 - 4b3=0

OpenStudy (anonymous):

so b2=10+20k a2 + b2= 0 so k=-5/9 now solve by your own..

OpenStudy (anonymous):

how did you get b2=10+20k...and thanks so much

OpenStudy (anonymous):

b1+b2+b3=5 +10k and -4 b1- 2b2 - 4b3=0 multiply 1st equation by 4 and add to 2nd equation...